OpenStudy (anonymous):

102+(-173)+266+(-185)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

102+(-173)=-71 -71+266=195 195+(-185)=10

OpenStudy (anonymous):

102 + (-173)=-71 266+(-185)=81 81+(-71)= 10 same answer different way to work.